OpenStudy (anonymous):

The function g is defined as @ganeshie8

OpenStudy (anonymous):

\[g:x \rightarrow \frac{ ax+b }{ x },x \neq~k\]

OpenStudy (anonymous):

Given that g(2)=3 and g(-2)=5,

OpenStudy (anonymous):

find the values of a and of b

ganeshie8 (ganeshie8):

\[g(x) = \dfrac{ax+b}{x}\]

ganeshie8 (ganeshie8):

se the given points and setup two equations

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ 2a+b }{ 2 }=3,\frac{ b-2a }{ -2 }=5\]

ganeshie8 (ganeshie8):

looks good! two equations and two unknowns, you can solve them

OpenStudy (anonymous):

\[2a+b=6,-2a+b=-10\]

OpenStudy (anonymous):

a=4,b=-2

OpenStudy (anonymous):

Thnx @ganeshie8

ganeshie8 (ganeshie8):

yw!

OpenStudy (anonymous):

:)